OpenStudy (anonymous):

use the definition to the derivative of : f(z)=(2z-i)/(z-2i) at z=-i .

OpenStudy (jamesj):

So, what's the problem here. The definition of derivative here is the limit as z --> i of \[\frac{f(z) - f(-i)}{z-(-i)}\] So write down that ratio and start simplifying until you can take the limit.

OpenStudy (anonymous):

the problem is that i get stuck half way before i can apply the limit ..

OpenStudy (phi):

One way to go is define z= -i + delta, and let delta approach zero: \[\lim_{\Delta \rightarrow 0}\frac{f(-i+\Delta)-f(-i)}{(-i+\Delta)-(-i)}\] This turns into \[\lim_{\Delta \rightarrow 0}\frac{1}{\Delta}\left\{ \frac{2(\Delta-i)-i}{(\Delta-3i)} -\left( \frac{-2i-i}{-3i} \right)\right\}\] Using a common denominator of \[-3i(\Delta-3i)\] add the two fractions, divide out terms, and then let Delta go to zero